a freight train leaves tamaqua, traveling east at 35km/hr. One hour later a passenger train leaves Tamaqua,also traveling east on a parallel track at 40km/hr. How far from tamaqua will the passenger train catch the freight train ?
Answer by stanbon(75887)

a freight train leaves tamaqua, traveling east at 35km/hr. One hour later a passenger train leaves Tamaqua,also traveling east on a parallel track at 40km/hr. How far from tamaqua will the passenger train catch the freight train ?
---------------
Freight Train DATA:
rate = 35 km/h ; time = x hrs ; distance = 35x km
--------------------------
Passenger Train DATA:
rate = 40 km/h ; time = x-1 hrs ; distance = 40(x-1) km
-----
Equation:
distance = distance
35x = 40x-40
-5x = -40
x = 8 hrs.
----
Distance from Tamaqua: 8*35 = 280 km
=========================================
Cheers,
Stan H.